OpenStudy (anonymous):

The u_rms of H2 molecules at 273K is 1.84*10^3 m/s At what temperature is u_rms for H2 1.9 times greater than this value?

OpenStudy (anonymous):

RMS velocity=sqrt(3RT/M). Apply the formula and get the answer.

OpenStudy (jfraser):

\[V{_R}{_M}{_S} = \sqrt{\frac{3RT}{MM}}\] It's asking for the temp at which the speed is roughly twice as fast (1.9x). Calculate the new V, rearrange and solve for the new T.